OpenStudy (anonymous):

How long will money in savings take to double at 5% interest compounded annually? Round the answer to the nearest hundredth of a year. Solve this application using logarithms.

ganeshie8 (ganeshie8):

Lets say the initial money = \(\large x\)

ganeshie8 (ganeshie8):

after doubling it would be : \(\large 2x\), and lets say it takes \(\large t\) years to double

ganeshie8 (ganeshie8):

use the compound interest formula : \(\large 2x = x(1+0.05)^t\) \(\large 2 = (1.05)^t\)

ganeshie8 (ganeshie8):

take log both sides and solve \(\large t\), can u ?

OpenStudy (anonymous):

I'm comfused